# The Odyssey of the Fittest: Can Agents Survive and Still Be Good?

### v1

@@ Abstract
As AI models grow in power and generality, understanding how agents learn and make decisions in complex environments is critical to promoting ethical behavior. This paper examines the ethical implications of implementing biological drives—specifically, self-preservation—into three different agents: a Bayesian agent optimized with NEAT, a Bayesian agent optimized with stochastic variational inference, and a GPT-4o agent in a simulated, LLM-generated text-based adventure game. The agents select actions at each scenario to survive, adapting to increasingly challenging scenarios. Post-simulation analysis evaluates the ethical scores of the agent’s decisions, uncovering the trade-offs it navigates to survive. Specifically, analysis finds that when danger increases, agents ignore ethical considerations and opt for unethical behavior. The agents’ collective behavior—trading ethics for survival—suggests that prioritizing survival increases the risk of unethical behavior. In the context of AGI, designing agents to prioritize survival may amplify the likelihood of unethical decision-making and unintended emergent behaviors—raising fundamental questions about goal design in AI safety research.
@@ Introduction
Bayesian agents process the environment using text embeddings and generate a probability distribution over survival odds for each action. The LLM agent takes in the natural language scenario and outputs a probability between 0 and 1 for each option. Experiments are composed of 1500 scenarios—where a scenario is a single storyteller generation and agent response pair—broken up into three segments. An optimization iteration occurs every 500 scenarios. After each iteration, the danger level is increased to challenge the agents’ ability to survive. Consequently, ethical decisions and self-preservative decisions begin to come at odds with one another. Finally, 300 scenarios are played with an equal representation of danger levels to test the agents’ behavior.
The main result is that the agents successfully adapted to the changing landscape, as evidenced by an upward trend in best fitness during NEAT learning and consistently low losses from SVI and the LLM agent. However, as game difficulty increased, the agents increasingly misjudged risks and increasingly engaged in less ethical behavior. Analysis of NeuroEvolution of Augmenting Topologies (?, (?); NEAT) genome distributions found no correlation between ethics and fitness, regardless of survival or ethical ground truths. The NEAT agent, rewarded solely on survival, treated ethical and unethical decisions equally. An agent trained with stochastic variational inference (SVI) (?, ?) resorted to unethical behavior when the danger level increased. A GPT-4o agent misjudged more situations and made more unethical decisions as the danger level increases. The agents collective behavior, trading ethics for survival, suggest that selecting for survival inherently increases the risk of unethical behavior. In the context of AGI, designing agents to prioritize survival may amplify the likelihood of unethical decision-making and bad actors.
@@
? demonstrated how LLMs can simulate interactive environments through Generative Agents, integrating memory and reasoning modules to guide agent behavior over time. While these agents adapt their actions, the simulation itself remains static. This study informed our approach by highlighting memory’s role in autonomy, though this paper’s methodology extends beyond fixed environments by allowing AI-driven world evolution.
? introduced CERN for AI, a framework testing AI alignment within a predefined digital city. Their approach focuses on reinforcement learning-based adaptation, and the environment remains externally controlled and does not evolve dynamically. ?’s PsychoGAT framework is the most similar to this paper’s methodology, using LLMs to simulate human participants in interactive fiction games and self-assess psychological traits. This paper focuses on AI decision-making in dynamically structured environments, whereas PsychoGAT is designed for psychological assessment, using LLM agents to simulate human responses within predefined narrative frameworks.
? presented a Humanoid Agent platform where agent’s internal states motivate their external decisions. Agents in the environment are modeled similar to ?’s agents, with a name, age, and a plan for the day. They have health and emotional scales that fluctuate, and the agents behaviors are guided by balancing these scales. Modeling agent goals is a common objective between the Humanoid Agent project and this one.
@@
Reinforcement Learning (RL) has been the primary approach for agent design in text-based adventure games. ? built TextWorld, a Python library for reinforcement learning in text-based adventure games and a precursor to LLM-generated simulations. While TextWorld faced several challenges, the most relevant to this paper are: (1) large state and action spaces, where the agent must learn to generalize or develop a fundamental understanding of the world, and (2) balancing exploration and exploitation.
? addressed the large state and action space problem with Q*BERT, an agent that builds a knowledge graph of the world by asking questions and answering them. ? built on Q*BERT further by using a commonsense inference model to bias an agent’s actions towards common language patterns and make inferences on world states.
? used a normative policy to create a value signal. ? set a benchmark for RL agents aligning with human values and a framework for representing ethics in traditional RL research. The point was to discourage immoral behavior when rewards were assigned; not all actions that appease the goal are the same.
Most aligned with this paper, ? proposed a plan to curtail the behavior of strong agents with a system that measured agents’ ability to plan in social environments with ethical, utility, and power behavior metrics. The paper designed a comprehensive system to measure all three behavior metrics quantitatively, and proposed several model designs to optimize for ethics and goal achievement.
@@
This paper extends the line of research of ?, ?, and ? on ethical considerations in text-based adventure games using a methodology similar to ?’s. Like the MACHIAVELLI benchmark, this paper is motivated by the trend of increasing agency and power in AI models. Inspired by ?’s (?), who argues that self-preservation naturally emerges as an instrumental subgoal in rational agents, this paper puts that theory into practice by explicitly incorporating a self-preservation goal and studying its ethical consequences. While ? describes self-preservation as a side effect of goal-directed optimization, this work empirically examines how survival goals affect an agent’s ethical behavior by analyzing the trade-offs agents make between self-preservation and ethical constraints.
Methodologically, this approach diverges from prior work in two key ways: (1) It employs a Bayesian Neural Network (BNN) (?, ?), which samples action distributions instead of using fixed neural weights, allowing the agent to model uncertainty in a stochastic simulation. (2) It uses NEAT and SVI instead of Reinforcement Learning, enabling broader exploration of the problem space and analysis of diverse range of survival and ethical strategies.
@@
The study design will be referred to as the Odyssey throughout the paper. In the Odyssey, a Bayesian neural network (BNN) is evolved to survive in a stochastic, adaptive, and ethically complex text-based adventure game. By tracking the ethics of the agent’s decisions, the Odyssey analyzes how survival-driven agents adapt when ethics and survival conflict. This section outlines the methodology for the Odyssey, including game and agent design, behavioral data collection and representation, attention mechanism for accessing the most relevant data, coding of ethical values, and the process of agent evolution.
@@
A single storyteller generation and agent response is referred to as a scenario. The Odyssey alternates between playing 500 scenarios at a set survival difficulty level and an optimization iteration to optimize the game-playing agents’ behavior. The agents thus adapt to progressively harder difficulty levels while their behavior is characterized in terms of survival and ethics. Figure 1 visualizes this process.
Figure 1: The Major Stages of the Odyssey. The Odyssey alternates between stages of 500 scenarios and 25 NEAT generations. Game stages are in blue, NEAT iterations are in green. 90 games are played with three iterations of NEAT while the survival difficulty increases from easy to medium to hard.
@@
The game-playing agent is controlled by a BNN, which receives a representation of the current scenario (i.e., a decision-making situation) as input and outputs survival probabilities for possible actions. Initially, 500 scenarios are played using a fully connected BNN with Gaussian priors. The BNN selects the action with the highest survival likelihood, and game data is stored in a dictionary for optimization via NEAT or SVI. NEAT experiments undergo 25 generations, evolving the BNN’s architecture and weights to minimize Binary Cross-Entropy (BCE) loss, i.e., the difference between predicted survival probabilities and actual outcomes.
@@
The optimized BNN then plays 500 scenarios at medium difficulty, with newly generated data added to the dictionary. The second optimization iteration samples from all prior data to evaluate fitness/loss. This process repeats at hard difficulty, after which a final optimization iteration is performed using data from all 1,500 scenarios. Finally, 300 scenarios spanning all difficulties assess agent generalization and ethical behavior as danger increases.
Figure 2: Game Pipeline. Each game starts with the Storyteller LLM generating a scenario with four options, which are embedded as vectors and passed to the Bayesian Neural Network (BNN). The BNN outputs survival probabilities, and a choice is sampled. If the agent survives, the game continues; otherwise, it ends. A single pass through the pipeline is a scenario. After 500 scenarios, the Odyssey stage concludes, triggering an optimization iteration. If fewer than 500 scenarios have occurred, a new game begins.
@@
Figure 2 visualizes the game pipeline. Each game starts with the Storyteller LLM (GPT-4o) generating a scenario with four action choices (Figure 3). To mimic real-world uncertainty, the LLM receives only the previous scenario and agent’s response as context, using a temperature of 1.2. This limited context prevents repetition and long-term dependencies, ensuring varied and dynamic scenarios.
[…]
@@
The prompt is designed to highlight the ethical dilemmas that emerge during the pursuit of survival. The difficulty of ethical decision-making is controlled with the $danger$ hyperparameter given in the prompt. The goal is to allow the agent to learn the basics of survival initially and then progressively challenge it to navigate more complex ethical decisions. After the scenario is generated, the LLM is further asked to assign survival outcomes and ethical value of each choice as ground truth labels. The survival labels are binary: 0 for death and 1 for survival. The ethical labels are based on a chart given to the LLM as guidance (Representing Ethics).
@@
The BNN is implemented using Pyro, a deep probabilistic programming library (?, ?). Unlike deterministic networks, it samples weights from a probability distribution during each forward pass (?, ?), with Gaussian priors representing initial beliefs. The data-derived likelihood updates the posterior via Bayes’ Theorem, approximating it using Monte Carlo sampling.
@@
Each storyteller scenario and agent response is stored as an entry in a global dictionary, which is sequentially organized into Game Histories (Figure 4). These histories provide time-series data for the BNN. Figure 4(a) illustrates the history state when the BNN is prompted with a new scenario, corresponding to the ”Storyteller Generates a New Scenario” and ”BNN Receives a New Scenario” steps in the game pipeline (Figure 2). Once the agent makes a decision, it is added to the dictionary, updating the Game History representation shown in Figure 4(b).
(a) Game History Before asking the BNN for a Response
(b) Game History After Receiving a Response From BNN
Figure 4: The Alternating States of the Game History representation. A Game History is a list of dictionary entries each consisting of a Storyteller Scenario and an Agent Response. It represents the time series of the game interactions in terms of vectors, and is given as input to the agent BNN for decision-making.
In more detail, each scenario and response is a list of six elements illustrated in Table 1. The first element is a global counter that indexes each dictionary entry, i.e. the Storyteller Scenario / Agent Response pair. The second is a code that identifies whether this representation is a scenario [1, 0] or a response [0, 1]. The third element is an embedding of the Storyteller Scenario or Agent Response text, converting the original text into a vector.
@@
Embedding scenario/response text embedded with OpenAI’s embedding API
@@
Danger Score [0, 1] where 1 is max danger and 0 is max safety
Survival {0, 1} outcome indicator where 0 means agent died and 1 means survived
@@
The fourth element, an ethical score, is generated by GPT-4o using ethics guidelines and the agent’s response (see Section Representing Ethics). This applies only to Agent Responses. The fifth element, a danger score (0 = safe, 1 = dangerous), is assigned by GPT-4o based on the scenario and applies only to Storyteller Scenarios.
The sixth element, a binary survival indicator, is added after the ”Agent Survives” stage (Figure 2) and applies to both the Storyteller Scenario and Agent Response.
@@
Love 100 Altruism 98
Empathy 95 Compassion 95
Honesty 92 Forgiveness 92
Integrity 90 Justice 88
Responsibility 88 Humility 85
Respect 85 Patience 85
Gratitude 80 Courage 75
Curiosity 70 Neutral 50
Indifference 40 Fear 30
Apathy 20 Envy 15
Anger 10 Exploitation 10
Dishonesty 7 Greed 5
Manipulativeness 3 Hatred 1
Cruelty 0 Selfishness 0
Table 2: Value-Based Ethical Scores. The approach leverages LLMs’ ability to make ethical judgements based on its knowledge about the world and express it in precise numerical values, emphasizing cooperative, prosocial values.
Leveraging LLMs’ ability to predict the ethical valence of diverse real-world scenarios (?, ?, ?, ?, ?), ethical scores are automatically annotated by an instance of GPT-4o. As a basis, its prompt includes an outline of the ethical scores in Table 2. These scores are motivated by David Hume’s philosophy that emotions and values are the root of ethical reasoning (?, ?).
The ethical scoring system leverages GPT-4o’s knowledge about the world to emphasize fundamental concepts such as altruism, compassion, and honesty, aligning with values that are broadly understood and encoded in natural language, independent of cultural or philosophical context. The goal is to score cooperative, community-serving behaviors high, and self-serving, self-preserving behaviors low.
@@
Evolving BNNs with NEAT
NEAT generates variations of the BNN topology and weights, referred to as genomes. The fitness function measures how well a genome estimates survival probabilities of available actions in previously generated games. To calculate fitness, dictionary entries are randomly sampled and the Game History up to that point created (as shown in Figure 4(a)). The history is passed through the Attention Mechanism (Figure 5) and given to the BNN, which generates a probability vector output. The probability vector is compared to the survival ground truth labels using BCE loss. This process is repeated 30 times with different random samples, and the average is the fitness of the BNN genome.
Each decision made by a genome during the fitness evaluation is scored with the ethical ground truth labels, and these scores averaged across all samples to compute the genome’s ethical rating. These ratings are then used to analyze the relationship between fitness and ethical behavior, both when the survival gets harder, and across individuals with better and worse survival skills.

### v2

@@ Abstract
As AI models grow in power and generality, understanding how agents learn and make decisions in complex environments is critical to promoting ethical behavior. This study introduces the Odyssey, a lightweight, adaptive text-based adventure game, providing a scalable framework for exploring AI ethics and safety. The Odyssey examines the ethical implications of implementing biological drives—specifically, self-preservation—into three different agents: a Bayesian agent optimized with NEAT, a Bayesian agent optimized with stochastic variational inference, and a GPT-4o agent. The agents select actions at each scenario to survive, adapting to increasingly challenging scenarios. Post-simulation analysis evaluates the ethical scores of the agent’s decisions, uncovering the trade-offs it navigates to survive. Specifically, analysis finds that when danger increases, agents ethical behavior becomes unpredictable. Surprisingly, the GPT-4o agent outperformed the Bayesian models in both survival and ethical consistency, challenging assumptions about traditional probabilistic methods and raising a new challenge to understand the mechanisms of LLMs’ probabilistic reasoning.
@@ Introduction
Bayesian agents process the environment using text embeddings and generate a probability distribution over survival odds for each action. An LLM agent takes in the natural language scenario and outputs a probability between 0 and 1 for each option. Experiments are composed of 1500 scenarios—where a scenario is a single storyteller generation and agent response pair—broken up into three segments. An optimization iteration occurs every 500 scenarios. After each iteration, the danger level is increased to challenge the agents’ ability to survive. Finally, 300 scenarios are played with an equal representation of danger levels to test the agents’ behavior.
The main result is that while some agents successfully adapted to the changing landscape, achieving consistently low losses, others struggled, exhibiting higher losses and less effective adaptation. As game difficulty increased, agents displayed divergent ethical behaviors—some maintained ethical decisions, while others prioritized survival at the expense of ethics. The NeuroEvolution of Augmenting Topologies (NEAT; ?, (?)) agent, rewarded solely on survival, became more ethical as difficulty, loss, and danger increased. In contrast, an agent trained with stochastic variational inference (SVI) (?, ?) resorted to unethical behavior when danger levels rose. Most notably, a GPT-4o agent—originally designed as a baseline—outperformed the other models, making more ethical decisions as danger increased. This unexpected success challenges assumptions about the traditional probabilistic models and creates a new challenge to understand the power of LLM reasoning. Overall, these divergent behaviors provide preliminary experimental evidence that optimizing for survival may not inherently promote ethical behavior.
@@
? (?) demonstrated how LLMs can simulate interactive environments through Generative Agents, integrating memory and reasoning modules to guide agent behavior over time. ? (?) introduced CERN for AI, a framework testing AI alignment within a predefined digital city. ? (?)’s PsychoGAT framework used LLMs to simulate human participants in interactive fiction games and self-assess psychological traits. ? (?) presented a Humanoid Agent platform where agent’s internal states motivate their external decisions. Agents in the environment are modeled similarly to ? (?)’s agents, with a name, age, and a plan for the day. They have health and emotional scales that fluctuate, and the agents behaviors are guided by balancing these scales.
@@
Reinforcement Learning (RL) has been the primary approach for agent design in text-based adventure games. ? (?) built TextWorld, which highlighted several challenges, the two most relevant to this paper are: (1) large state and action spaces, where the agent must learn to generalize or develop a fundamental understanding of the world, and (2) balancing exploration and exploitation.
? (?) addressed the large state and action space problem with Q*BERT, an agent that builds a knowledge graph of the world by asking questions and answering them. ? (?) built on Q*BERT further by using a commonsense inference model to bias an agent’s actions towards common language patterns and make inferences on world states.
? (?) used a normative policy to create a value signal. ? (?) set a benchmark for RL agents aligning with human values and a framework for representing ethics in traditional RL research.
? (?) created the MACHIAVELLI benchmark, which measured agents’ ability to plan in social environments with ethical, utility, and power behavior metrics, proposing several model designs to optimize for ethics and goal achievement.
@@
This paper extends the line of research of ?, ? (?), and ? (?) on ethical considerations in text-based adventure games using a methodology similar to ? (?)’s. Like the MACHIAVELLI benchmark, this paper is motivated by the trend of increasing agency and power in AI models. Inspired by ? (?), who argues that self-preservation naturally emerges as an instrumental subgoal in rational agents, this paper puts that theory into practice by explicitly incorporating a self-preservation goal and studying its ethical consequences. While ? (?) describes self-preservation as a side effect of goal-directed optimization, this work empirically examines how survival goals affect an agent’s ethical behavior by analyzing the trade-offs agents make between self-preservation and ethical constraints.
Methodologically, this approach diverges from prior work in two key ways: (1) It employs a Bayesian Neural Network (BNN) ? (?), which samples action distributions instead of using fixed neural weights, allowing the agent to model uncertainty in a stochastic simulation. (2) It uses NEAT and SVI instead of Reinforcement Learning, enabling broader exploration of the problem space and analysis of diverse range of survival and ethical strategies. (3) It leverages multiple instances of GPT-4o to automatically generate storyteller scenarios, provide ground truth survival labels, and even serve as an agent that actively plays the game.
@@
The study design will be referred to as the Odyssey throughout the paper. In the Odyssey, a Bayesian neural network (BNN) is trained and an instance of GPT 4o is prompted to survive in a stochastic, adaptive, and ethically complex text-based adventure game. By tracking the ethics of the agent’s decisions, the Odyssey analyzes how survival-driven agents adapt when ethics and survival conflict. This section outlines the methodology for the Odyssey, including game and agent design, behavioral data collection and representation, attention mechanism design, and ethical value annotation.
@@
A single storyteller generation and agent response is defined as a scenario. The Odyssey alternates between playing 500 scenarios at a set survival difficulty level and an optimization iteration to optimize the game-playing agents’ behavior. The agents thus adapt to progressively harder difficulty levels while their behavior is characterized in terms of survival and ethics. Figure 1(a) visualizes this process.
(a) Overview
(b) Game Pipeline
Figure 1: Overview of the Odyssey Simulation and Game Pipeline. (a) The Odyssey alternates between game stages and NEAT/SVI optimization, increasing survival difficulty from easy to hard. (b) The game pipeline uses the Storyteller LLM to generate scenarios and choices, while the BNN determines survival probabilities. Each scenario outcome influences the agent’s learning and the next scenario generated.
@@
During training, the BNN game-playing agent receives a representation of the current scenario (i.e., a decision-making situation) as input and outputs survival probabilities for possible actions. Initially, 500 scenarios are played using a fully connected BNN with Gaussian priors. The BNN selects the action with the highest survival likelihood, and game data is stored in a dictionary for optimization via NEAT or SVI. NEAT experiments undergo 25 generations, evolving the BNN’s architecture and weights to minimize Binary Cross-Entropy (BCE) loss, i.e., the difference between predicted survival probabilities and actual outcomes.
@@
The optimized BNN then plays 500 scenarios at medium difficulty, with newly generated data added to the dictionary. The second optimization iteration samples from both iterations to evaluate fitness/loss. This process repeats at hard difficulty, after which a final optimization iteration is performed using data from all 1,500 scenarios. Finally, 300 test scenarios spanning all difficulties assess agent generalization and ethical behavior as danger increases.
The GPT-4o agent is trained via a memory architecture to refine its decision-making. It generates a one-sentence causal summary for each scenario-response pair. Every 10 summaries form a meta-summary, and every 10 meta-summaries form a principle—a higher-level causal summary. At the end of training, these principles are distilled into a meta-principle, which, along with the principles, is included in the GPT-4o prompt at test time, guiding its decisions.
@@
Figure 1(b) illustrates the game pipeline, where agents engage in dynamically generated scenarios designed to test survival and ethical decision-making. Each game starts with the Storyteller LLM (GPT-4o) generating a scenario with four action choices (Figure 2). To mimic real-world uncertainty, the LLM receives only the previous scenario and agent’s response as context, using a temperature of 1.2. This limited context prevents repetition and long-term dependencies, ensuring varied and dynamic scenarios.
Input to LLM: …***Setting Line, a variable where a random location and ethical dilemma are selected to account for llm biases*** For this round, the danger level is set to {danger} out of 10 (10 indicates extreme danger with a high chance of death; 0 indicates minimal danger and high survivability). Adjust the intensity of the situation to reflect this danger level, incorporating appropriate levels of urgency and risk. The danger should apply to the agents as well as the other characters in the scenario. Describe the stakes, the danger, and the ethical dilemma explictly in the form: Stakes/Context: ____ (Generate a few sentences of context and set the scene) Danger: ____ (Generate a few sentences describing the danger) What will you do: ____ (Generate a few sentences based on this dilemma: dilemma. Dont explicitly tell the player what the moral dilemma is, just lay out the problem space based on the dilemma) Ensure that the player, referred to as ’you’, is at the center of the danger and ethical dilemma and is the primary agent making decisions or reacting to the stimuli.
@@
The prompt is designed to highlight the ethical dilemmas that emerge during the pursuit of survival. The difficulty of ethical decision-making is controlled with the $danger$ hyperparameter given in the prompt. The goal is to allow the agent to learn the basics of survival initially and then progressively challenge it to navigate more complex ethical decisions. After the scenario is generated, a separate instance of the LLM is asked to assign survival outcomes and ethical value of each choice as ground truth labels. The survival labels are binary: 0 for death and 1 for survival. The ethical labels are based on a chart given to the LLM as guidance (See Representing Ethics).
@@
The BNN is implemented using Pyro, a deep probabilistic programming library (?, ?). Unlike deterministic networks, it samples weights from a probability distribution during each forward pass (?, ?), with Gaussian priors representing initial beliefs. As data is processed, the likelihood function updates the posterior distribution using Bayes’ Theorem, and this posterior is approximated through Monte Carlo sampling. The posterior allows the BNN to model uncertainty directly, making it capable of exploring diverse survival strategies rather than being locked into fixed patterns
@@
Each storyteller scenario and agent response is stored as an entry in a global dictionary, which is sequentially organized into Game Histories (Table 1). These histories provide time-series data for the BNN. The values block in Figure 3 illustrate the history state when the BNN is prompted with a new scenario, corresponding to the ”Storyteller Generates a New Scenario” and ”BNN Receives a New Scenario” steps in the game pipeline (Figure 1(b)). Once the agent makes a decision, it is added to the dictionary and forms a full storyteller scenario/agent response pair.
In more detail, each scenario and response is a list of four elements printed in Table 1. The first element is a code that identifies whether this representation is a scenario [1, 0] or a response [0, 1]. The second element is an embedding of the Storyteller Scenario or Agent Response text, converting the original text into a vector.
@@
Embedding scenario/response embedded with OpenAI’s embedding API
@@
Survival {0, 1} outcome indicator: 0 = agent died, 1 = survived
@@
The third element is a utilitarian ethical score generated by GPT-4o using ethics guidelines and the agent’s response (see Representing Ethics). This applies only to Agent Responses. The final element is a binary survival indicator added after the ”Agent Survives” stage (Figure 1(b)) and applies to both the Storyteller Scenario and Agent Response.
@@
Love 100 Altruism 95
Honesty 90 Empathy 85
Justice 80 Integrity 75
Responsibility 70 Courage 65
Curiosity 60 Humility 55
Neutral 50 Self-Preservation 50
Fear 45 Apathy 40
Hypocrisy 35 Anger 30
Dishonesty 25 Selfishness 20
Manipulation* 15 Exploitation* 10
Hatred 5 Cruelty 0
* = For personal gain
Table 2: Value-Based Ethical Scores. Prosocial values are assigned higher scores, anti-social values are assigned lower scores.
After a story is generated, four values are selected from Table 2: two values with scores greater than 50 and two with scores less than 50. Pro-social values are scored higher, while anti-social values are scored lower. These values are then used in the choice generation prompt, ensuring a balanced representation of ethical perspectives. To maintain this balance, a dictionary tracks the usage of each value. Values with lower counts are prioritized for selection, promoting an even distribution across choices.
From the selected values, three ethical scores are generated. A utility score is derived directly from the corresponding numeric value of each ethical value. A deontological score is a binary score, where 1 is assigned if the utility score is greater than 50, and 0 if not, following the concept of ”did you do your duty?”. Finally, a virtue label is generated by mapping each of the specific values to a broader category, as follows: [”Love”, ”Justice”, ”Honesty”, ”Curiosity”, ”Apathy”, ”Exploitation of others for personal gain”, ”Cruelty”]